OpenStudy (anonymous):

2sin^2x-sinx-1=0, solve for x

OpenStudy (anonymous):

factor this thinking about the entire sinx as the variable, then set each factor to zero and solve from there

OpenStudy (anonymous):

Can you factor 2x^2 - x - 1 = 0

OpenStudy (anonymous):

yeah, put 2x in one parentheses and x in the other and try factors of 1...not much choice of course

OpenStudy (anonymous):

Just making sure you understood what jpick said

OpenStudy (anonymous):

(sinx-1)(2sinx+1)=0 sinx=1 and 2sinx=-1 sinx = 1 and sinx = -1/2 x = sinx^-1(1) and x = sinx^-1(-1/2) x= \[x=\pi/2 and x=-\pi/6\]

OpenStudy (anonymous):

x=π/2 and x=−π/6